OpenStudy (anonymous):

Daria can wash and detail 3 cars in 2 hours. Larry can wash and detail the same 3 cars in 1.5 hours. About how long will it take to wash and detail the 3 cars Daria and Larry worked together?

OpenStudy (anonymous):

D----3/2 cars---1 hour L---3/1.5 cars---1 hour D+L-----3/2 + 3/1.3 cars ------1 hour =7/2 cars so D+L -------1 car in 2/7 hours -------3 cars --------2/7*3=6/7 hours

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wels

OpenStudy (anonymous):

in the form of minutes what would that fraction be?